OpenStudy (anonymous):

you have been hired to make a roller coaster more exciting. the owners want the speed at the bottom of the first hill doubled. how much higher must the first hill be built?

OpenStudy (jamesj):

What's the formula for kinetic energy? It is \[ KE = \frac{1}{2}mv^2 \] If we double the speed at the bottom from v to 2v, the KE changes to KE' \[ KE \ ' = \frac{1}{2}m(2v)^2 = 2mv^2 = 4KE \] In other words, the kinetic energy increases by a factor 4. Now, that being the case, how much higher does the hill have to be?

OpenStudy (anonymous):

As when there is no change in velocity it's kinetic energy is 1/2mv^2.where v is velocity and m is mass when velocity is doubled then kinetic energy will be 4 times using above formula.i.e kinetic energy increases four times to previous value...and calculate what you want...

OpenStudy (anonymous):

is the answer four times?